OpenStudy (anonymous):

solve 3x^2-7x-5=0 and find x intercepts

OpenStudy (saifoo.khan):

the answer will be in radicals.

OpenStudy (saifoo.khan):

\[\frac 16 (7- \sqrt{109}),\frac 16 (7+ \sqrt{109})\]

OpenStudy (anonymous):

thanks, i have a few more to post and then i need to eat, so will be back in about an hour or so for appendix f if you will be on then

OpenStudy (saifoo.khan):

Alright! :)